OpenStudy (anonymous):

whrite 3 ratos equal to 4/28

OpenStudy (mr.math):

\[\frac{4}{28}=\frac{2}{14}=\frac{1}{7}\]

OpenStudy (nubeer):

2:14 1:7 8:56